OpenStudy (anonymous):

Let R and S be rings and ϕ:R→S be an isomorphism, the ϕ is __________________ monomorphism endomorphism Automorphism homomorphism

OpenStudy (loser66):

homo for sure

OpenStudy (anonymous):

An example of an endomorphic map is ___________________ ϕ:Z→Q ϕ:R→C ϕ:Z→Z ϕ:C→R

OpenStudy (anonymous):

i think c @Loser66

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Let R be the ring[0,1]of all conditions real valued function on the close interval [0,1] and ϕ:R→R defined by ϕ(f)=(1/2)f then ϕ is _______ homomorphism Endomorphism Isomorphism Epimorphism

OpenStudy (anonymous):

@Loser66 are you sure or you think?

OpenStudy (loser66):

now, i am sure, since before you typed 12 f, it is nonsense

OpenStudy (loser66):

now it is 1/2 f, makes more sense

OpenStudy (anonymous):

have a lot more please help

OpenStudy (anonymous):

If G is agroup and H is a subgroup of G.then the order of G divides the order H the order of H divides the order G H1∩G2 has same order as H H1∩G2 has same order as G

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

the order of H divides the order of G

OpenStudy (loser66):

I don't know what is H1 or G2

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Consider the subset I={ 0,3,6} of the ring equivalent of [0] is ------------------ [8,10,12,14,...] [...,-18,-9,0,9,18,...] [...,30,36,42,48,54,..] [...,2,4,6,8,10,...]

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

What is the ring?

OpenStudy (anonymous):

z9

OpenStudy (loser66):

set of multiple of 9

OpenStudy (loser66):

You should understand what is going on

OpenStudy (loser66):

in \(\mathbb Z_9\) the congruence [0] is set of number divided by 9

OpenStudy (loser66):

for example 9 |9 =1 Remainder 0 , this remainder is what it is indicate to congruence 0

OpenStudy (loser66):

9|18 =2 Remainder 0, hence 18 is in the set you are looking for

OpenStudy (loser66):

Suppose I give you 20, then 9|20 = 2 Remainder 2, hence 20 is NOT in the set of congruence 0

OpenStudy (loser66):

Among the options, only {....., -18,-9,0,9,18.......} all terms are multiple of 9. Each term divided by 9 get 0 in remainder. Hence this set is the answer. Got me?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Let ϕ:R→R′ be a homomorphism of a ring R into a ring R'. The one of these is not true ϕ(0)=0′ ϕ(−a)=−ϕ(a) ϕ(e)=e′ ϕ′(s′) is not a subring of R where S' is a subring

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

what is \(\phi '\)

OpenStudy (loser66):

im back in 30 minutes

OpenStudy (anonymous):

ok sir. will be waiting

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you there?

OpenStudy (loser66):

yes

OpenStudy (anonymous):

ok are you back?

OpenStudy (loser66):

but I don't know what is phi ' from the last one, is it inverse of phi?

OpenStudy (anonymous):

dont know . dats just the question

OpenStudy (anonymous):

i think it is inverse

OpenStudy (loser66):

Actually, among them, the last one seems not true since the homomorphism doesn't give us that homomorphism is onto (surjective) , hence nothing guarantee that S' has pre-image in R

OpenStudy (anonymous):

ok

OpenStudy (loser66):

Moreover, all the other choices are correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let R be the ring[0,1]of all conditions real valued function on the close interval [0,1] and ϕ:R→R defined by ϕ(f)=1/2f what is Kerφ

OpenStudy (loser66):

0

OpenStudy (loser66):

let f is in kernel, then \(\phi (f) = 1/2 f =0\) if and only if f =0, Hence f =0 is kernel phi

OpenStudy (anonymous):

Let G be a group and H1 , H2 normal subgroups of .one of these is a normal subgroup of G H1∩H2 H1∪H2 H1−H2 A×B

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):